If you exert 12.6 N for a distance of 2.4 m in 1.5 s, you will deliver a power of __W, N
1
Expert's answer
2019-07-24T15:49:35-0400
"\rm{Power}=\frac{{Work\:done}}{\rm{Time}}"

"P=\frac{W}{t}=\frac{F\times d}{t}"

"=\frac{12.6\:\rm{N}\times 2.4\:\rm{m}}{1.5\:\rm{s}}=20.16\:\rm{W}"